# Echoes and defects in the Calogero model

### v1

@@ IV Validation through experimentally relevant protocols
In the following, we study two experimentally-relevant protocols beyond the sudden quenches studied in the literature Rajabpour and Sotiriadis (2014); del Campo (2016). To explore the inherently out-of-equilibrium and non-adiabatic regimes, we consider the Floquet protocol
$\omega^{2}(t)\equiv\omega^{2}_{0}(1+h\cos(\Omega t)),$ (9)
@@
with rate $\delta$ and $t\in[-1/\delta,1/\delta]$. In other words, the gas is slowly let free of its trap before being trapped again. The quantity $z\nu$ in Eq. (10) is a priori simply a parameter describing the dynamical protocol. The model may however be viewed as an effective model for a many-body system ramped across its quantum critical point in the spirit of Refs. Defenu et al. (2018b); Polkovnikov and Gritsev (2008); Bachmann et al. (2017); Defenu (2021). Within this perspective the exponent $z\nu$ may represent the dynamical critical exponent for the gap scaling. The point at which the trap is zero corresponds to the gapless critical point in this system, where the energy levels are degenerate. This specific functional form of the protocol is of particular interest because of its universality in the adiabatic limit: any drive through $\omega=0$, provided it is slow enough, can be expanded to lowest order yielding the form given by Eq. (10). We therefore expect the response of the system to this protocol to encompass the response of all drives which can be expanded into a power law in the adiabatic limit. The two protocols and their effects on the many-body spectrum are illustrated in Fig. 2.
@@ V Results
The Ermakov-Milne equation (6) can be solved for both protocols: in terms of non-linear combinations of Mathieu functions for the Floquet drive in Eqs. (9) and generalized Airy functions for the slow drive protocol in (10). The details of the calculations in the case of the Floquet drive are reproduced in Appendix B, while the corresponding computations for the drive through the gapless point closely follow those obtained in Defenu (2021). Given Eq. (7), this yields an exact expression for the evolved ground state. As detailed in Appendix A in agreement with del Campo (2016), the ground-state fidelity can be derived in terms of the solution to the Ermakov-Milne equation and reads
@@
We first address the case of the Floquet protocol. This study is of great interest given that the underlying $SO(2,1)$ symmetry structure of the Calogero Hamiltonian is intimately linked to the emergence of parametric resonance. It is well known that parametric instabilities can emerge in such systems when both the amplitude $h$ and the modulation frequency $\Omega$ are varied. In the $(h,\tfrac{\Omega}{\omega_{0}})$ space, parametric resonance delineates regions which are stable from those which are unstable giving rise to a pattern of stability lobes (known as Arnold tongues Nayfeh and Mook (1995)). The stability Gritsev and Polkovnikov (2017) and periodicity Lv et al. (2020) properties of the system’s response to the drive (9) are universal features associated with the $SO(2,1)$ dynamical symmetry. In particular, the occurrence of particular beat structures and parametric instabilities is uniquely determined by the characteristic exponent of the Mathieu equation associated with the drive (9), which depends only on $h$ and the ratio $\omega_{0}/\Omega$. Note that the parametric instability corresponds to the points in $(h,\omega_{0}/\Omega)$-space where the Mathieu exponent becomes complex.
@@
Perfect ground state revivals on the other hand are directly linked to the commensurability of the two frequencies $\omega_{0}$ and $\Omega$. For all $\lambda$ and $N$, these perfect revivals occur at times which are solely determined by the parameters $(h,\omega_{0}/\Omega)$. These times are shared identically by the classical Mathieu oscillator, parametric quantum harmonic oscillator (corresponding to $N=1,\lambda=0$) and the corresponding periodically-driven trapped free Bose ($N>1,\lambda=0$) and Tonks-Girardeau gas ($N>1,\lambda=1$), with no effect of interactions $\lambda$. Given the $\lambda$-independence of these features, the periodically-driven Calgero model (which generalizes the latter models to $N>1,\lambda>0$) will display analog beat structures, perfect echoes and instabilities dictated uniquely and solely by parameters of the drive $(h,\omega_{0}/\Omega)$. The universal stability diagram Gritsev and Polkovnikov (2017) associated with periodically-driven systems with $SO(2,1)$ dynamical symmetry (including the Calogero model and the other systems mentioned above) is displayed in Fig. 3.
@@
While the occurrence of perfect echoes and the beat structures are solely dictated by $(h,\omega_{0}/\Omega)$, we will see that the echo amplitude in between exact revival points is greatly influenced by the interactions $\lambda$, consistently with Eq. (15). In other words, $\lambda$ crucially determines the likeliness of echoes beyond those ensured by commensurability of the drive and the natural trapping frequency. This reveals an intricate and strong interdependence between interactions/statistics and echoes in the Floquet setting.
@@
Our results for the ground state echo amplitude $e(t)$ of the Calogero model in the parametrically stable regimes of the Floquet drive are displayed in Fig. 4 for fixed $N$ and several interaction strengths $\lambda$, and this for three sets of parameters $(\tfrac{\omega_{0}}{\Omega},h)$ in the stable region with different degrees of stability. The position of the corresponding regimes in parameter space are indicated in the stability diagram displayed in Fig. 3.
@@
Figure 3: Stability diagram of periodically-driven systems with $SO(2,1)$ dynamical symmetry. Inset: zoom on the parametric values considered in Fig. 4. These correspond to regimes where the echo amplitude is (a) dominated by the natural trap frequency $\omega_{0}$ for low drive amplitudes $h$, (b) a result of the competition between the driving frequency of the trap $\Omega$ and $\omega_{0}$ for intermediate drive amplitudes $h$, (c) dominated by the driving frequency of the trap $\Omega$ for drive amplitudes $h$ approaching the parametric instability.
@@
Given commensurate frequencies ${2\omega_{0}}$ and ${\Omega}$, one observes that at times where the two corresponding sinusoidal signals (say $\cos(2\omega_{0})$ and $\cos(\Omega)$) are in phase, an echo occurs with absolute certainty: the corresponding amplitude is unity and is unaffected by $\lambda$. In contrast, at times where the former are not in phase, the echo amplitude is strongly suppressed by larger values of $\lambda$ through the exponential suppression embodying Eq. (15). This is most visible in the intermediate regime of the drive’s amplitude $h$, when both the natural and drive frequencies compete. Here, one witnesses a dramatic suppression of the intermediate peaks occurring between the commensurability-imposed perfect echoes (see Fig. 4b). In the small $h$ regime, the echo amplitude is dominated by the natural frequency of the trap $2\omega_{0}$ as seen in Fig. 4a, while for large values of $h$ (approaching the parametric instability), the echo amplitude is dominated by the frequency of the drive $\Omega$ as seen in Fig. 4c. The effective frequency structure in intermediate regimes such as those of Fig. 4b is dictated by the Mathieu exponent, which depends only on the parameters of the drive (and not $\lambda$), and interpolates between those two cases. In the limiting cases of vanishing $h$ or amplitudes $h$ approaching the instability, intermediate echoes would become ever more likely and the resulting perfect echoes would occur with frequencies $2\omega_{0}$ and $\Omega$ respectively. Figs. 4a and 4c however clearly demonstrate that small deviations from these limits already spark large drops in the likelihood of intermediate echoes. To summarize, we observe under Floquet drives a universal decrease in the chances of intermediate revivals through interactions and exclusion (beyond those perfect echoes imposed by commensurability).
@@
Figure 4: Ground state echo amplitude $e(t)$ for a periodic drive with a non-resonant but commensurate drive frequency $\Omega=3/2$ in units of the trap frequency $\omega_{0}$, for $N=10$. Each subpanel displays $e(t)$ for an amplitude of the drive $h$ given by (a) 1/8 (b) 1/3 (c) 1/2. The curves in each panel from top to bottom correspond to $\lambda=0,\tfrac{1}{4},\tfrac{1}{2},1,\tfrac{3}{2},2$. The points considered in parameter space of the drive are referenced within the stability digram in Fig. 3.
@@
Figure 5: Ground state fidelity $f(t)$ for a slow quench through $\omega=0$ with parameters $z\nu\equiv 1/4,\;N=10$ and from top to bottom $\lambda\in\{0,\tfrac{1}{4},\tfrac{1}{2},1,\tfrac{3}{2},2\}$.
@@
$\lim_{t\rightarrow\infty}f(t)=(\sin(p\pi))^{N(1+\lambda(N-1))},$ (16)
with $p=(2+2z\nu)^{-1}$, showing that the asymptotic fidelity of a harmonic quantum Calogero gas driven across its critical point is always a constant irrespectively of the rate exponent $z\nu$ of the ramp time dependence. While this constant depends on $z\nu$ as well as $N$ and $\lambda$, it is universal in its independence from the rate $\delta$. This is consistent with the universal breakdown of of adiabaticity predicted to follow the crossing of infinitely degenerate quantum critical points (encountered in effective free bosonic theories) Defenu (2021); Bachmann et al. (2017). Our results however extend those predictions to interacting settings or equivalently to non-trivial exclusion statistics: the asympotic values (16) display a clear dependence on $\lambda$. As seen in Fig. 5, interactions (or exclusion statistics) $\lambda$ induce a further drop in the asymptotic fidelity. This is consistent with the enhanced proliferation and hindered annihilation of defects through interactions predicted in section III.
@@ VI Discussion
In summary, we have investigated the interplay between exclusion statistics and spectrum generating symmetries in the harmonically confined Calogero model out of equilibrium. Our studies considered both periodic and adiabatic modulations of the trap potential, focusing on the ground-state echo amplitudes and fidelity. Both quantities are dramatically suppressed with increasing exclusion quantified by the interaction strength $\lambda$. Given a periodic drive and a commensurate natural trap frequency, we predict a sharp suppression of the likelihood of intermediate echoes (between those imposed by commensurability) in the presence of interactions. Additionally, our study of the effects of interactions forecasts a substantial drop in the asymptotic ground state fidelity reached following a slow drive through $\omega=0$, independent of the driving rate $\delta$. All of these findings are consistent with favored proliferation and hindered annihilation of defects sparked by interactions, which can be interpreted as a result of exclusion statistics.
These conclusions are underpinned by the intricate and fascinating interplay between statistics and dynamical symmetries of scale-invariant fluids embodied by the Calogero model, which we hope to verify in near-term future experiments. Following this study, it is of great interest to look into further observables such as the pair-momentum distribution. Although perhaps less easily amenable to analytical treatment, the latter is highly relevant in cold atom experiments and would provide great additional insights regarding the consequences of statistics and scaling dynamics on correlations and spatial coherence. Further work of great significance includes the investigation of dynamics of the Calogero model breaking the $SO(2,1)$ dynamical symmetry through modulation through time of the interaction parameter $\lambda$ into a regime where quantum anomalies and bound states may play a role or the study of multispecies and spinful generalizations of the Calogero model. Our work also motivates the exploration of the interplay between fractional statistics and dynamics in two dimensional systems like quantum Hall systems hosting anyons.
@@ Acknowledgements.
This work was supported by a collaborative ETH Grant, the Swiss National Science Foundation (SNSF) under project funding ID: 200021_207537, by the Swiss Secretariat for Education, Research and Innovation (SERI) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster).
Appendix A Derivation of ground state fidelity and echo amplitudes
Given the dynamical state’s evolution (7), we now derive expressions for observable quantities of particular interest. In particular, we first focus on the ground state echo amplitude and fidelity, given respectively by the overlap of the dynamical state (evolved from the initial state taken as the ground state) and the instantaneous equilibrium (or adiabatic) ground state. The latter is given by
@@ Appendix A Derivation of ground state fidelity and echo amplitudes
This is obtained simply by simply replacing $\omega$ in the expression of the static ground state of the Calgero model (2) by its dynamic equivalent $\omega(t)$. The ground state fidelity then refers to
@@
where the implied domain of integration is ${\mathbb{R}^{N}}$. By inserting the expressions (7) and (17) above for the dynamical and adiabatic ground states, we find that
@@
with $\Omega(t)=-i\frac{\dot{\xi}(t)}{\xi(t)}+\frac{\bar{\omega}}{\xi^{2}(t)}$. To obtain a closed form for the expression above, we make use of Mehta’s integral Met (2004)
@@
$=(2\pi)^{N/2}\prod_{j=0}^{N-1}\frac{\Gamma(1+(j+1)\lambda)}{\Gamma(1+\lambda)}=\frac{\mathcal{N}_{N,\lambda}}{2^{-\frac{N}{2}[1+\lambda(N-1)]}}.$
@@
is found by rescaling $x_{i}$ by a factor $|\omega(t)+\Omega(t)|$. This respectively yields factors $|\omega(t)+\Omega(t)|^{N}$ from the measure and $|(\omega(t)+\Omega(t))|^{\lambda N(N-1)}$ from the Jastrow factor. Inserting this result and the value of Metha’s integral, the following closed expression for the ground state fidelity is obtained
$f(t)=\left(\frac{4\omega(t)\bar{\omega}}{\alpha(t)\xi^{2}(t)}\right)^{N[1+\lambda(N-1)]/2}$ (22)
@@
$\alpha(t)=\left[\left(\omega(t)+\frac{\bar{\omega}}{\xi(t)^{2}}\right)^{2}+\left(\frac{\dot{\xi}(t)}{\xi(t)}\right)^{2}\right]$ (23)
@@
which is obtained by replacing $\omega(t)$ by $\bar{\omega}=\omega(t_{0})$ in the expression of the overlap with the adiabatic equilibrium ground state

### v2

@@ IV Validation through experimentally relevant protocols
In the following, we study two experimentally-relevant protocols beyond the sudden and linear quenches studied in the literature Rajabpour and Sotiriadis (2014); del Campo (2016, 2021). To explore the inherently out-of-equilibrium and non-adiabatic regimes, we consider the Floquet protocol
$\omega^{2}(t)\equiv\omega^{2}_{0}[1+h\cos(\Omega t)],$ (9)
@@
with rate $\delta$ and $t\in[-1/\delta,1/\delta]$. In other words, the gas is slowly let free of its trap before being trapped again. The quantity $z\nu$ in Eq. (10) is a priori simply a parameter describing the dynamical protocol. The model may, however, be viewed as an effective model for a many-body system ramped across its quantum critical point in the spirit of Refs. Defenu et al. (2018a); Polkovnikov and Gritsev (2008); Bachmann et al. (2017); Defenu (2021). Within this perspective the exponent $z\nu$ may represent the dynamical critical exponent for the gap scaling. The point at which the trap is zero corresponds to the gapless critical point in this system, where the energy levels are degenerate. This specific functional form of the protocol is of particular interest because of its universality in the adiabatic limit: any drive through $\omega=0$, provided it is slow enough, can be expanded to lowest order yielding the form given by Eq. (10). We, therefore, expect the response of the system to this protocol to encompass the response of all drives which can be expanded into a power law in the adiabatic limit. The protocols and their effects on the many-body spectrum are illustrated respectively in Figs. 2 and 3.
Figure 3: Schematic representation of the Floquet experimental protocol investigated in this work. Starting with the system in its ground-state (in gray in the leftmost black box) trapped with a given frequency $\omega=\omega_{+}$, the protocol drives the trapping frequency to a lower value $\omega_{-}$ before being restored, this procedure being repeated with a period $2\pi/\Omega$.
@@ V Results
The Ermakov-Milne equation (6) can be solved for both protocols: in terms of non-linear combinations of Mathieu functions for the Floquet drive in Eqs. (9) and generalized Airy functions for the slow drive protocol in Eq.(10). The details of the calculations in the case of the Floquet drive are reproduced in Appendix B, while the corresponding computations for the drive through the gapless point closely follow those obtained in Ref. Defenu (2021). Given Eq. (7), this yields an exact expression for the evolved ground-state. As detailed in Appendix A in agreement with Ref. del Campo (2016), the ground-state fidelity can be derived in terms of the solution to the Ermakov-Milne equation and reads
@@
We first address the case of the Floquet protocol. This study is of great interest given that the underlying $\mathrm{SO(2,1)}$ symmetry structure of the Calogero Hamiltonian is intimately linked to the emergence of parametric resonance. It is well known that parametric instabilities can emerge in such systems when both the amplitude $h$ and the modulation frequency $\Omega$ are varied. In the $(h,\tfrac{\Omega}{\omega_{0}})$ space, parametric resonance delineates regions which are stable from those which are unstable giving rise to a pattern of stability lobes (known as Arnold tongues Nayfeh and Mook (1995)). The stability Gritsev and Polkovnikov (2017) and periodicity Lv et al. (2020) properties of the system’s response to the drive (9) are universal features associated with the $\mathrm{SO(2,1)}$ dynamical symmetry. In particular, the occurrence of particular beat structures and parametric instabilities is uniquely determined by the characteristic exponent of the Mathieu equation associated with the drive (9), which depends only on $h$ and the ratio $\omega_{0}/\Omega$. Note that the parametric instability corresponds to the points in $(h,\omega_{0}/\Omega)$-space where the Mathieu exponent becomes complex.
@@
Perfect ground-state revivals, however, are directly linked to the commensurability of the two frequencies $\omega_{0}$ and $\Omega$. For all $\lambda$ and $N$, these perfect revivals occur at times which are solely determined by the parameters $(h,\omega_{0}/\Omega)$. These times are shared identically by the classical Mathieu oscillator, parametric quantum harmonic oscillator (corresponding to $N=1,\lambda=0$) and the corresponding periodically driven trapped free Bose ($N>1,\lambda=0$) and Tonks-Girardeau gas ($N>1,\lambda=1$), with no effect of interactions $\lambda$. Given the $\lambda$-independence of these features, the periodically driven Calgero model (which generalizes the latter models to $N>1,\lambda>0$) will display analog beat structures, perfect echoes and instabilities dictated uniquely and solely by parameters of the drive $(h,\omega_{0}/\Omega)$. The universal stability diagram Gritsev and Polkovnikov (2017) associated with periodically driven systems with $\mathrm{SO(2,1)}$ dynamical symmetry (including the Calogero model and the other systems mentioned above) is displayed in Fig. 4.
@@
While the occurrence of perfect echoes and the beat structures are solely dictated by $(h,\omega_{0}/\Omega)$, we will see that the echo amplitude in between exact revival points is greatly influenced by the interactions $\lambda$, consistently with Eq. (15). In other words, $\lambda$ crucially determines the likeliness of echoes beyond those ensured by commensurability of the drive and the natural trapping frequency. This reveals an intricate and strong interdependence between interactions and statistics and echoes in the Floquet setting.
@@
Our results for the ground-state echo amplitude $e(t)$ of the Calogero model in the parametrically stable regimes of the Floquet drive are displayed in Fig. 5 for fixed $N$ and several interaction strengths $\lambda$, and this for three sets of parameters $(\tfrac{\omega_{0}}{\Omega},h)$ in the stable region with different degrees of stability. The position of the corresponding regimes in parameter space are indicated in the stability diagram displayed in Fig. 4.
@@
Figure 4: Stability diagram of periodically driven systems with $\mathrm{SO(2,1)}$ dynamical symmetry. Inset: Zoom-in on the parametric values considered in Fig. 5. These correspond to regimes where the echo amplitude is (a) dominated by the natural trap frequency $\omega_{0}$ for low drive amplitudes $h$, (b) a result of the competition between the driving frequency of the trap $\Omega$ and $\omega_{0}$ for intermediate drive amplitudes $h$, and (c) dominated by the driving frequency of the trap $\Omega$ for drive amplitudes $h$ approaching the parametric instability.
@@
Given commensurate frequencies ${2\omega_{0}}$ and ${\Omega}$, one observes that at times where the two corresponding sinusoidal signals [say $\cos(2\omega_{0})$ and $\cos(\Omega)$] are in phase, an echo occurs with absolute certainty: the corresponding amplitude is unity and is unaffected by $\lambda$. In contrast, at times where the former are not in phase, the echo amplitude is strongly suppressed by larger values of $\lambda$ through the exponential suppression embodying Eq. (15). This is most visible in the intermediate regime of the drive’s amplitude $h$, when both the natural and drive frequencies compete. Here, one witnesses a dramatic suppression of the intermediate peaks occurring between the commensurability-imposed perfect echoes [see Fig. 5(b)]. In the small $h$ regime, the echo amplitude is dominated by the natural frequency of the trap $2\omega_{0}$ as seen in Fig. 5(a), while for large values of $h$ (approaching the parametric instability), the echo amplitude is dominated by the frequency of the drive $\Omega$ as seen in Fig. 5(c). The effective frequency structure in intermediate regimes such as those of Fig. 5(b) is dictated by the Mathieu exponent, which depends only on the parameters of the drive (and not $\lambda$), and interpolates between those two cases. In the limiting cases of vanishing $h$ or amplitudes $h$ approaching the instability, intermediate echoes would become ever more likely and the resulting perfect echoes would occur with frequencies $2\omega_{0}$ and $\Omega$ respectively. Figs. 5(a) and 5(c) however clearly demonstrate that small deviations from these limits already spark large drops in the likelihood of intermediate echoes. To summarize, we observe under Floquet drives a universal decrease in the chances of intermediate revivals through interactions and exclusion (beyond those perfect echoes imposed by commensurability).
@@
Figure 5: ground-state echo amplitude $e(t)$ for a periodic drive with a non-resonant but commensurate drive frequency $\Omega=3/2$ in units of the trap frequency $\omega_{0}$, for $N=10$. Each subpanel displays $e(t)$ for an amplitude of the drive $h$ given by (a) 1/8 (b) 1/3, and (c) 1/2. The curves in each panel from top to bottom correspond to $\lambda=0,\tfrac{1}{4},\tfrac{1}{2},1,\tfrac{3}{2},2$. The points considered in parameter space of the drive are referenced within the stability digram in Fig. 4.
@@
Figure 6: ground-state fidelity $f(t)$ for a slow quench through $\omega=0$ with parameters $z\nu\equiv 1/4,\;N=10$ and from top to bottom $\lambda\in\{0,\tfrac{1}{4},\tfrac{1}{2},1,\tfrac{3}{2},2\}$.
@@
$\lim_{t\rightarrow\infty}f(t)=[\sin(p\pi)]^{N(1+\lambda(N-1))},$ (16)
with $p=(2+2z\nu)^{-1}$, showing that the asymptotic fidelity of a harmonic quantum Calogero gas driven across its critical point is always a constant irrespectively of the rate exponent $z\nu$ of the ramp time dependence. While this constant depends on $z\nu$ as well as $N$ and $\lambda$, it is universal in its independence from the rate $\delta$. This is consistent with the universal breakdown of of adiabaticity predicted to follow the crossing of infinitely degenerate quantum critical points (encountered in effective free bosonic theories) Defenu (2021); Bachmann et al. (2017). Our results, however, extend those predictions to interacting settings or equivalently to non-trivial exclusion statistics: the asympotic values (16) display a clear dependence on $\lambda$. As seen in Fig. 6, interactions (or exclusion statistics) $\lambda$ induce a further drop in the asymptotic fidelity. This is consistent with the enhanced proliferation and hindered annihilation of defects through interactions predicted in Sec. III.
@@ VI Discussion
In summary, we investigated the interplay between exclusion statistics and spectrum generating symmetries in the harmonically confined Calogero model out of equilibrium. Our studies considered both periodic and adiabatic modulations of the trap potential, focusing on the ground-state echo amplitudes and fidelity. Both quantities are dramatically suppressed with increasing exclusion quantified by the interaction strength $\lambda$. Given a periodic drive and a commensurate natural trap frequency, we predict a sharp suppression of the likelihood of intermediate echoes (between those imposed by commensurability) in the presence of interactions. Additionally, our study of the effects of interactions forecasts a substantial drop in the asymptotic ground-state fidelity reached following a slow drive through $\omega=0$, independent of the driving rate $\delta$. All of these findings are consistent with favored proliferation and hindered annihilation of defects sparked by interactions, which can be interpreted as a result of exclusion statistics.
These conclusions are underpinned by the intricate and fascinating interplay between statistics and dynamical symmetries of scale-invariant fluids embodied by the Calogero model, which we hope to verify in near-term future experiments.
Although $1/r^{2}$ interactions have not yet been the subject of targeted experiments, a possible realization of the Calogero model in cold Bose atom systems with dipole-dipole interactions was devised in Ref. Yu (2006). Meanwhile one also could potentially engineer a many-body system out of Efimov states Macek (2007), which are bound in an effective $1/r^{2}$ potential or exploit the tunability of interactions offered through the use of dipolar molecular gases or mediation via fermions or guided photons Argüello-Luengo et al. (2022); Hung et al. (2016); Li et al. (2021). A direct observation of defects in the fashion of Ref. Ko et al. (2019) would then allow for experimental observation of the proliferation of defects with enhanced statistics predicted in our work. An even cleverer approach would be to construct and study in a similar fashion SO(2,1) symmetric systems tailored to model interactions already realized in the existing experimental platforms.
Following this study, it is also of great interest to look into further observables such as the pair-momentum distribution. Although perhaps less easily amenable to analytical treatment, the latter is highly relevant in cold atom experiments and would provide great additional insights regarding the consequences of statistics and scaling dynamics on correlations and spatial coherence. Further work of great significance includes the investigation of dynamics of the Calogero model breaking the $\mathrm{SO(2,1)}$ dynamical symmetry through modulation through time of the interaction parameter $\lambda$ into a regime where quantum anomalies and bound states may play a role or the study of multispecies and spinful generalizations of the Calogero model. Our work also motivates the exploration of the interplay between fractional statistics and dynamics in two dimensional systems like quantum Hall systems hosting anions.
@@ Acknowledgements.
This work was supported by an ETH Zürich Research Grant, the Swiss National Science Foundation (SNSF) under project funding ID: 200021_207537, by the Swiss Secretariat for Education, Research and Innovation (SERI) and by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC2181/1-390900948 (the Heidelberg STRUCTURES Excellence Cluster).
Appendix A Derivation of ground-state fidelity and echo amplitudes
Given the dynamical state’s evolution (7), we now derive expressions for observable quantities of particular interest. In particular, we first focus on the ground-state echo amplitude and fidelity, given respectively by the overlap of the dynamical state (evolved from the initial state taken as the ground-state) and the instantaneous equilibrium (or adiabatic) ground-state. The latter is given by
@@ Appendix A Derivation of ground-state fidelity and echo amplitudes
This is obtained simply by simply replacing $\omega$ in the expression of the static ground-state of the Calgero model (2) by its dynamic equivalent $\omega(t)$. The ground-state fidelity then refers to
@@
where the implied domain of integration is ${\mathbb{R}^{N}}$. By inserting the expressions (7) and (17) above for the dynamical and adiabatic ground-states, we find that
@@
with $\Omega(t)=-i\frac{\dot{\xi}(t)}{\xi(t)}+\frac{\bar{\omega}}{\xi^{2}(t)}$. To obtain a closed form for the expression above, we make use of Mehta’s integral Mehta (2000)
@@
$=(2\pi)^{N/2}\prod_{j=0}^{N-1}\frac{\Gamma[1+(j+1)\lambda]}{\Gamma(1+\lambda)}=\frac{\mathcal{N}_{N,\lambda}}{2^{-\frac{N}{2}[1+\lambda(N-1)]}}.$
@@
is found by rescaling $x_{i}$ by a factor $|\omega(t)+\Omega(t)|$. This respectively yields factors $|\omega(t)+\Omega(t)|^{N}$ from the measure and $|[\omega(t)+\Omega(t)]|^{\lambda N(N-1)}$ from the Jastrow factor. Inserting this result and the value of Metha’s integral, the following closed expression for the ground-state fidelity is obtained
@@
$\alpha(t)=\left[\left(\omega(t)+\frac{\bar{\omega}}{\xi(t)^{2}}\right)^{2}+\left(\frac{\dot{\xi}(t)}{\xi(t)}\right)^{2}\right].$ (23)
@@
which is obtained by replacing $\omega(t)$ by $\bar{\omega}=\omega(t_{0})$ in the expression of the overlap with the adiabatic equilibrium ground-state